OpenStudy (anonymous):

Find the slope of the line passing through the points (3, 4) and (2, 1). Option A: 1 Option B:- 1 Option C: 3/5 Option D: -5/3

OpenStudy (anonymous):

Find Slope. (3, 4) and (2, 1) (x1, y1) and (x2, y2) \[slope = \frac{ y2 - y1 }{ x2 - x1 }\] \[slope = \frac{ 1 - 4 }{ 2 - 3 } = \frac{ -3 }{ -1 } = 3\] You have a slope of 3.